The average time for a courier to travel from Manila to Pampanga is 200 minutes, and the standard deviation is 10 minutes. If one of these trips is selected at random, find the probability that a courier will have the following travel time: a. At least 180 minutes b. At most 205 minutes
1
Expert's answer
2020-10-29T20:52:04-0400

Assume that the travel times are normally distributed.

a. "P(X\ge180)"

"\mu=200,\sigma=10"

"P(X\ge180)=P(Z\ge\frac{180-200}{10})"

"=P(Z\ge-2)"

"=1-P(Z\le-2)"

"=1-0.02275"

"=0.97725"

b. "P(X\le205)"

"P(X\le205)=P(Z\le\frac{205-200}{10})"

"=P(Z\le0.5)=0.6915"